What is the surface area of regular pentagonal pyramid with base area of 49 square meters and slant height of 13 meters.
Answer by addingup(3677)

See attached illustration to learn how we found the length of the sides of the base (5 equal sides).
OK, so you looked at the illustration and now have the formulas.
As you see, 5.34 is the length of the sides of the base. Now looking up we have 5 triangles each with a base 5.34 and a height of 13.
The formula for the area of the triangle is A= bh/2:
A= 5.34*13/2
A= 34.71 is the area of each triangular side. And we have 5 sides:
34.71*5= 173.55 Add area of base:
173.55+49 222.55 is the surface area of this pentagon.